OpenStudy (anonymous):

an entry test conducted demanded that the test takers answers 45 quantative questions within 55 minutes. which of the following is closest to the average amount of time the taker can spend on each question?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

1.2222222 min

OpenStudy (anonymous):

if 3x+2y=1 & 2x+3y=4 then x+y must be?

OpenStudy (anonymous):

@mrchen

OpenStudy (anonymous):

if 3x+2y=1 & 2x+3y=4 then x+y must be? @freckles

OpenStudy (anonymous):

3 i think

OpenStudy (anonymous):

eh prob not.